OpenStudy (some.random.cool.kid):

Solve the quadratic equation x2 − 6x + 13 = 0 using the quadratic formula. What is the solution when expressed in the form a ± bi, where a and b are real numbers? x = 3 ± 2i x = 2 ± 3i x = 3 ± 4i x equals start fraction three plus or minus two i over two end fraction

OpenStudy (some.random.cool.kid):

please walk me thorugh...

OpenStudy (some.random.cool.kid):

through* to get the right answr. @Nnesha

OpenStudy (some.random.cool.kid):

@peachpi

OpenStudy (some.random.cool.kid):

@pooja195

OpenStudy (some.random.cool.kid):

@DanJS @Data_LG2 @dan815 @triciaal

OpenStudy (danjs):

you know the quadratic formula that results from solving that standrard quadratic form for x =

OpenStudy (some.random.cool.kid):

i knew it but forgot mind showing me it?

OpenStudy (danjs):

If ax^2 + bx + c = 0 \[x = \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a}\]

OpenStudy (some.random.cool.kid):

yes yess ik that ah thank ok so b is b 13 is c and 1 is a correct?

OpenStudy (some.random.cool.kid):

thank you*

OpenStudy (some.random.cool.kid):

oops i said b is bxD

OpenStudy (some.random.cool.kid):

i meant b is 6

OpenStudy (danjs):

a = 1 b = -6 c = 13

OpenStudy (some.random.cool.kid):

oops -6 not 6 ok

OpenStudy (danjs):

x^2 + (-6)x .....

OpenStudy (some.random.cool.kid):

uh solve that?

OpenStudy (some.random.cool.kid):

want me to solve it?

OpenStudy (danjs):

yeah , use those values and simplify as much as you can

OpenStudy (some.random.cool.kid):

uh...wait step by step please what will this look like when finished finding each letter ot the number? can you rewrite for me but this time instead of variables substitute it with numbers so that i may have an idea in my mind i do better solving when i have a visual.

OpenStudy (danjs):

ill type it out fast, i have to run for a while ... one sec

OpenStudy (some.random.cool.kid):

ok thank you

OpenStudy (danjs):

the battery died in my mouse right after that comment, sorry took forever to find a new one

OpenStudy (danjs):

\[x = \frac{ - (-6) \pm \sqrt{(-6)^2-4*1*13} }{ 2*1 }\]

OpenStudy (danjs):

\[x = \frac{ 6 \pm \sqrt{-16} }{ 2 }\]

OpenStudy (danjs):

the square root of a negative is not in the real numbers, for the complex, remember that the imaginary number i = square root of -1

OpenStudy (some.random.cool.kid):

so so do i find the sqrt now because its 4

OpenStudy (danjs):

\[x = \frac{ 6 \pm \sqrt{16 * -1} }{ 2 } = \frac{ 6 \pm \sqrt{16}*\sqrt{-1} }{ 2 }\]

OpenStudy (danjs):

replace root of -1 with that imaginary I symbol , then just have to simplify by dividing 2 into the top , root 16 is 4

OpenStudy (some.random.cool.kid):

i did say 4 its above

OpenStudy (danjs):

\[x = \frac{ 6 \pm 4*i}{ 2 } = 3 \pm 2i\]

OpenStudy (some.random.cool.kid):

isnt that the simpliest of forms?

OpenStudy (some.random.cool.kid):

so quick question do you have time for two more?

OpenStudy (danjs):

yeah, that is a conjugate pair, complex numbers come in pairs like that ..+ and -

OpenStudy (danjs):

x = 3 + 2i and x = 3 - 2i

OpenStudy (some.random.cool.kid):

you've been the only great help today that was able to explain in manor that i understood hope you have more time for jus two more.

OpenStudy (danjs):

sure , i can spend like 10 min prolly

OpenStudy (some.random.cool.kid):

great that should be enough. next one says -

OpenStudy (some.random.cool.kid):

Solve the equation 4x2 + 8x + 1 = 0 by completing the square. x equals start fraction three plus or minus square root of six end square root over two end fraction x equals six plus or minus two square root of six end square root x equals start fraction negative two plus or minus square root of three end square root over two end fraction x equals start fraction four plus or minus three square root of six end square root over eight end fraction

OpenStudy (some.random.cool.kid):

is this same method or is this the other method with teh factoring?

OpenStudy (danjs):

completing square, you want to turn the ax^2 + bx + c into something that will factor into (term)*(term) or (term)^2, which are sides of equal length of a square..

OpenStudy (danjs):

as now you cant... 4x^2 + 8x + 1 = 0

OpenStudy (danjs):

you can add something to both sides of the equations so that you find a value that will factor 4x^2 + 8x + c The thing to remember is that you just take half of b^2

OpenStudy (danjs):

(b=8) [ b/2]^2 = [8/2]^2 = 16

OpenStudy (some.random.cool.kid):

so 8/2^2?

OpenStudy (some.random.cool.kid):

oh yeah

OpenStudy (danjs):

add 16 to both sides of the quadratic like this... (4x^2 + 8x + 16) + 1 = 0 + 16

OpenStudy (danjs):

factorable now, and simplify

OpenStudy (danjs):

4(x^2 + 2x + 4) = 16 - 1

OpenStudy (some.random.cool.kid):

mmm wait you had me and then lost me lol so what am i factoring the 16 8x and the other one if so how do i go about doing this, by finding the common multiple?

OpenStudy (danjs):

The reason to add (b/2)^2 = 16 to both sides, was because that will make the first two terms in the quadratic a perfect square with it... factors easy

OpenStudy (danjs):

see i added the 16 and grouped it with the 4x^2 + 8x + 16 ready to factor

OpenStudy (some.random.cool.kid):

alright so...

OpenStudy (danjs):

4(x^2 + 2x + 4) = 16 - 1

OpenStudy (danjs):

x^2 + 2x + 4 = 15/4 (x + 2)*(x + 2) = 15/4

OpenStudy (some.random.cool.kid):

15/4 = 3.75 correct?

OpenStudy (some.random.cool.kid):

still there?

OpenStudy (some.random.cool.kid):

mouse died again? lawl

OpenStudy (danjs):

no , messed up, answer was wrong... i see where now... before you complete the square by adding on that half b squared term, you have to simplify the 4x^2 + 8x terms first by taking the factor of 4 out...

OpenStudy (danjs):

i will type it all

OpenStudy (danjs):

4x^2 + 8x + 1 = 0 4*(x^2 + 2x) + 1 = 0 x^2 + 2x = -1/4 now complete square

OpenStudy (danjs):

x^2 + 2x + 1 = -1/4 + 1 (x + 1)^2 = 3/4

OpenStudy (some.random.cool.kid):

mmm ok

OpenStudy (some.random.cool.kid):

is there something i should do next?

OpenStudy (some.random.cool.kid):

dont leave ill be right back using the restroom real quick.

OpenStudy (danjs):

yeah take the square root of both sides, which recall has a + and a - value then subtract the 1 to isolate x

OpenStudy (danjs):

\[x+1 =\pm \sqrt{3/2} = \frac{ \pm \sqrt{3} }{ 2 }\] \[x = \frac{ \pm \sqrt{3} }{ 2 } - 1\]

OpenStudy (some.random.cool.kid):

ok

OpenStudy (some.random.cool.kid):

whats next?

OpenStudy (danjs):

that is the answer, i was looking at the choices... x equals start fraction negative two plus or minus square root of three end square root over two end fraction

OpenStudy (some.random.cool.kid):

this should end with -2 + or minus sqrt of 3 all over two correct?

OpenStudy (some.random.cool.kid):

oh nvm you just said that...

OpenStudy (danjs):

they wrote it as a single fraction, one last step i guess if you wanted.... replace the -1 with - 2/2

OpenStudy (danjs):

then combine into a single fraction

OpenStudy (some.random.cool.kid):

k got it :D this prob took longer than i thought sure you still have time for the last one?

OpenStudy (danjs):

cuz i f'ed up in like the first step.. hah

OpenStudy (some.random.cool.kid):

lol

OpenStudy (danjs):

remember to put the thing into simler form next time before anything else

OpenStudy (some.random.cool.kid):

ok

OpenStudy (danjs):

simplest form, standard form

OpenStudy (danjs):

sure , quickly,

OpenStudy (some.random.cool.kid):

k great

OpenStudy (some.random.cool.kid):

What are the zeros for the function f(x) = 3x2 + 2x + 2 and how many times does the graph cross the x-axis? x equals start fraction negative one plus or minus i square root of seven end square root over three end fraction comma one time x equals start fraction negative one plus or minus i square root of five end square root over three end fraction comma never x equals start fraction negative two plus or minus two i square root of five end square root over three end fraction comma two times x equals start fraction negative one plus or minus five i over three end fraction comma never

OpenStudy (some.random.cool.kid):

i believe i got this one i thnk its d but it could also be the other never..

OpenStudy (some.random.cool.kid):

my teach said it would be never but idk which one i figured the imaginary number but i might be wrong.

OpenStudy (danjs):

zeroes are x-intercept values, y = f(x) = 0 0 = 3x^2+2x+2

OpenStudy (some.random.cool.kid):

ok

OpenStudy (danjs):

it is the second one , x equals start fraction negative one plus or minus i square root of five end square root over three end fraction comma never

OpenStudy (some.random.cool.kid):

man so it was the second one darn knew it could figure out which one thanks i almost would have said the one because after i solved it i gave up on trying to figure out whether it was an imag or no...

OpenStudy (danjs):

that one you just hav eto put the numumbers into tha tquadratic formula for a = 3 b=2 c=2

OpenStudy (some.random.cool.kid):

thank you for your help and patience. night i am done with this review my next i dont think requires assistant so i will let you go now :~>

OpenStudy (danjs):

yeah no real zeros, just complex roots

OpenStudy (danjs):

complex zeros

OpenStudy (danjs):

k, i should be on here tomorrow , not sure

OpenStudy (some.random.cool.kid):

:D i pasted all this as review for my EOC at the end of this semester well midterm ...

OpenStudy (some.random.cool.kid):

ok thats perfect

OpenStudy (some.random.cool.kid):

do you have an idea of what time tomorrow i have some more stuff thats related to this...

OpenStudy (some.random.cool.kid):

i can hopefully use these as notes...